# Inflation via Moduli Potentials in a Nested Warped Geometry

### v1

@@ Abstract
We analyze the effective four-dimensional dynamics of the extra-dimensional moduli fields in curved braneworlds having nested warping, with particular emphasis on the doubly warped model which is interesting in the light of current collider constraints on the mass of the Kaluza-Klein graviton. The presence of a non-zero brane cosmological constant ($\Omega$) naturally induces an effective moduli potential in the four-dimensional action, which shows distinct features in dS ($\Omega>0$) and AdS ($\Omega<0$) branches. For the observationally interesting case of dS 4-branes, a metastable minimum in the potential arises along the first modulus, with no minima along the higher moduli. The underlying nested geometry also leads to interesting separable forms of the non-canonical kinetic terms in the Einstein frame, where the brane curvature directly impacts the kinetic properties of only the first modulus. We subsequently explore the ability of curved multiply warped geometries to drive inflation with an in-built exit mechanism, by assuming predominant slow roll along each modular direction on a case-by-case basis. We find slow roll on top of the metastable plateau along the first modular direction to be the most viable scenario, with the higher-dimensional moduli parametrically tuning the height of the potential without significant impact on the inflationary observables. On the other hand, while slow roll along the higher moduli can successfully inflate the background and eventually lead to an exit, consistency with observations seemingly requires unphysical hierarchies among the extra-dimensional radii, thus disfavouring such scenarios.
@@ 1 Introduction
Motivated by the reasons highlighted so far, we have generalized multiply warped braneworld scenarios to include non-zero brane curvature and explored their geometric properties in an earlier work [61]. In the present work, our aim is to study the resulting induced potential in multiply warped non-flat backgrounds, and to focus on some of its cosmological implications vis-à-vis the singly warped non-flat model. In Sec. 2, we start with a concise review of nested warped geometry in presence of non-zero brane curvature. We subsequently discuss the origin of the 4D induced potential in the curved doubly warped scenario in Sec. 3, and derive the effective non-canonical action for the moduli fields in the Einstein frame in terms of the Jordan frame action. In Sec. 4, we investigate the scope of the resulting 4D effective framework in driving slow roll inflation while maintaining consistency with current observations. Moving on to Sec. 5, we highlight a few dynamical features which appear generic to non-flat geometries with more than two levels of nested warping. We conclude by summarizing our key results and discussing a few possible future directions in Sec. 6.
@@ 2 Review of non-flat multiply warped scenario
where $a(y)$ and $b(z)$ are the two warp factors, and $T_{1}$ and $T_{2}$ are the orbifold radii. Note that the intrinsic 3-brane metric $g_{\mu\nu}$ is not identified a priori with $\eta_{\mu\nu}$, i.e. one allows room for the possibility that the “end of the world” branes may be curved. This curvature can be interpreted as the result of a 4D cosmological constant, $\Omega$, induced geometrically on the corner branes, that non-trivially modifies the forms of both the warp factors and the brane tensions compared to the flat scenario. There are two possible branches depending on whether $\Omega<0$ or $\Omega>0$, which correspond to anti-de Sitter (AdS) and de Sitter (dS) geometries respectively.
@@ 2.1 Anti-de Sitter ($\Omega<0$)
where $c$ and $k$ are the dimensionless moduli corresponding to the $y$ and $z$ orbifolds respectively. We define $\omega_{1}^{2}=-\dfrac{\Omega T_{1}^{2}}{3c^{2}}$ to be the dimensionless induced cosmological constant, and $c_{1}=1+\sqrt{1-\omega_{1}^{2}}$. The moduli and the radii are not all indepenent, but remain algebraically related via two equalities: $k=T_{2}\sqrt{-\dfrac{\Lambda_{6}}{40M^{4}}}$ and $c=\dfrac{T_{1}k}{T_{2}\textrm{cosh}(k\pi)}$. The 4-brane tensions, obtained by consistently matching boundary conditions, turn out to be
@@ 3.2 Effective kinetic and curvature terms
with $k^{\prime}=\sqrt{-\dfrac{\Lambda_{6}}{40M^{4}}}$ extracted to make the Jordan frame coupling $h(c,k)$ dimensionless. The explicit forms of $h(c,k)$ for the AdS and dS cases, obtained by plugging in the corresponding warp factors in the integral above, are given in (A.1) and (A.2) in Appendix A.
@@ 3.3 Transformation from Jordan to Einstein frame
Clearly, this action contains a non-minimal coupling of curvature to the moduli fields through $h(c,k)$. This coupling can be made minimal by transforming to the Einstein frame, which is enacted by first assuming
@@
Figure 1: Representative plots of the Einstein frame non-canonical kinetic coupling functions $\hat{\tilde{A}}_{kk}(k)$ on the upper left, $\hat{\tilde{A}}_{cc}(c)$ on the upper right, $\hat{\tilde{A}}_{ck}(c,k)$ at the lower left, and the Einstein frame induced potential $\hat{V}(c,k)$ at the lower right, for AdS geometry with $\omega_{1}=10^{-15}$. Note that $\hat{V}(c,k)$ is in units of $10^{-15}$.
@@
Figure 2: Representative plots of the Einstein frame non-canonical kinetic coupling functions $\hat{\tilde{A}}_{kk}(k)$ on the upper left, $\hat{\tilde{A}}_{cc}(c)$ on the upper right, $\hat{\tilde{A}}_{ck}(c,k)$ at the lower left, and the Einstein frame induced potential $\hat{V}(c,k)$ at the lower right, for dS geometry with $\omega_{2}=10^{-3}$. Note that $\hat{V}(c,k)$ is in units of $10^{-6}$.
@@
The kinetic coefficients and the induced potential in the Einstein frame have been plotted in figure 1 for the AdS case and in figure 2 for the dS case. The explicit forms of the AdS kinetic coefficients are given in Appendix A under (A.1), (A.1), and (A.1), and the dS ones under (A.2), (A.2), and (A.2). The corresponding functional forms of $\hat{V}(c,k)$ appear in (A.1) and (A.2). It is interesting to note that $\hat{\tilde{A}}_{cc}$ depends on $c$ only and $\hat{\tilde{A}}_{kk}$ is a function of $k$ alone, while $\hat{\tilde{A}}_{ck}$ is a function of both $c$ and $k$. In this sense, the Einstein frame allows a nice decomposition of the functional dependence of the coupling coefficients based on the kinetic terms they are associated with, which is a consequence of the underlying nested warped geometry and holds good in higher dimensional extensions. Also, $\hat{\tilde{A}}_{kk}(k)$ is independent of $\Omega$, and as such, is identical across the AdS and dS scenarios. These features appear to be generic properties which can be readily generalized to higher-dimensional extensions, as we discuss later in Sec. 5. On the other hand, $\hat{\tilde{A}}_{cc}(c)$ in the dS case changes sign from negative to positive (for increasing $c$) at a non-zero value of $c$, unlike in the AdS case where it approaches zero from the negative side asymptotically. This feature of the dS regime turns out to be crucial from a cosmological point of view, as we shall demonstrate in the upcoming section.
@@
A few words about $\hat{V}(c,k)$ in the de Sitter regime are in order. In the 5D curved scenario, it is known that the induced dS potential may stabilize the extra-dimensional modulus by virtue of its metastable minimum [58, 59]. While these earlier studies have been carried out in terms of the canonically normalized radion field derived from the modulus, the metastable nature of the plateau should reflect in the modular potential as well. In fact, if one proceeds directly in terms of the modulus in the 5D set-up, the shape of the induced potential that one ends up with is identical to a constant $k$-projection of $\hat{V}(c,k)$ from figure 2, as expected physically. Such a $k$-slice clearly shows a nearly flat plateau in the range $c\in[1,5]$. So here is the natural question to ask: might it be possible to stabilize the other moduli of multiply warped geometries in a similar fashion? Based on figure 2, we find the answer to be negative. In fact, $\hat{V}(c,k)$ shows no local minimum, metastable or not, in the direction of the second modulus $k$, along which it decays monotonically. As we observe later in Sec. 5, this trend continues to hold in nested warped extensions beyond 6D, with the induced potential displaying no local minimum along the higher-dimensional moduli in field space. Thus, in multiply warped extensions of the 5D curved scenario, one indeed appears to need additional phenomenological mechanisms and/or beyond-Einstein gravitational effects to stabilize the higher moduli beyond the first one.

### v2

# Moduli Dynamics in Effective Nested Warped Geometry in Four Dimensions and Some Cosmological Implications

@@ Abstract
We analyze the effective four-dimensional dynamics of the extra-dimensional moduli fields in curved braneworlds having nested warping, with particular emphasis on the doubly warped model which is interesting in the light of current collider constraints on the mass of the Kaluza-Klein graviton. The presence of a non-zero brane cosmological constant ($\Omega$) naturally induces an effective moduli potential in the four-dimensional action, which shows distinct features in dS ($\Omega>0$) and AdS ($\Omega<0$) branches. For the observationally interesting case of dS 4-branes, a metastable minimum in the potential arises along the first modulus, with no minima along the higher moduli. The underlying nested geometry also leads to interesting separable forms of the non-canonical kinetic terms in the Einstein frame, where the brane curvature directly impacts the kinetic properties of only the first modulus. The non-canonicity of the scenario has been illustrated via an explicit computation of the field space curvature. We subsequently explore the ability of curved multiply warped geometries to drive inflation with an in-built exit mechanism, by considering predominant slow roll along each modular direction on a case-by-case basis. We find slow roll on top of the metastable plateau along the first modular direction to be the most viable scenario, with the higher-dimensional moduli parametrically tuning the height of the potential without significant impact on the inflationary observables. On the other hand, while slow roll along the higher moduli can successfully inflate the background and eventually lead to an exit, consistency with observations seemingly requires unphysical hierarchies among the extra-dimensional radii, thus disfavouring such scenarios.
@@ 1 Introduction
Motivated by the reasons highlighted so far, we have generalized multiply warped braneworld scenarios to include non-zero brane curvature and explored their geometric properties in an earlier work [61]. In the present work, our aim is to study the resulting induced potential in multiply warped non-flat backgrounds, and to focus on some of its cosmological implications vis-à-vis the singly warped non-flat model. In Sec. 2, we start with a concise review of nested warped geometry in presence of non-zero brane curvature. We subsequently discuss the origin of the 4D induced potential in the curved doubly warped scenario in Sec. 3, derive the effective non-canonical action for the moduli fields in the Einstein frame in terms of the Jordan frame action, and demonstrate the non-canonicity of the scenario under consideration by explicitly computing the curvature of the field space metric. In Sec. 4, we investigate the scope of the resulting 4D effective framework in driving slow roll inflation while maintaining consistency with current observations. Moving on to Sec. 5, we highlight a few dynamical features which appear generic to non-flat geometries with more than two levels of nested warping. We conclude by summarizing our key results and discussing a few possible future directions in Sec. 6.
@@ 2 Review of non-flat multiply warped scenario
where $a(y)$ and $b(z)$ are the two warp factors, and $T_{1}$ and $T_{2}$ are the orbifold radii of mass dimension $-1$. Note that the intrinsic 3-brane metric $g_{\mu\nu}$ is not identified a priori with $\eta_{\mu\nu}$, i.e. one allows room for the possibility that the “end of the world” branes may be curved. This curvature can be interpreted as the result of a 4D cosmological constant, $\Omega$, induced geometrically on the corner branes, that non-trivially modifies the forms of both the warp factors and the brane tensions compared to the flat scenario. There are two possible branches depending on whether $\Omega<0$ or $\Omega>0$, which correspond to anti-de Sitter (AdS) and de Sitter (dS) geometries respectively.
@@ 2.1 Anti-de Sitter ($\Omega<0$)
where $c$ and $k$ are the dimensionless moduli corresponding to the two orbifolds, with $y$ and $z$ being their respective angular coordinates. We define $\omega_{1}^{2}=-\dfrac{\Omega T_{1}^{2}}{3c^{2}}$ to be the dimensionless induced cosmological constant, and $c_{1}=1+\sqrt{1-\omega_{1}^{2}}$. The moduli and the radii are not all indepenent, but remain algebraically related via two equalities: $k=T_{2}\sqrt{-\dfrac{\Lambda_{6}}{40M^{4}}}$ and $c=\dfrac{T_{1}k}{T_{2}\textrm{cosh}(k\pi)}$. The 4-brane tensions, obtained by consistently matching boundary conditions, turn out to be
@@
Finally, an interesting interpretation of the 6D model can be developed based on the fundamental brane tensions. Recall that the tension of the 4-brane situated at $z=0$ vanishes, i.e. $V_{3}(y)=0$, in both the AdS and the dS branches, as noted respectively in (2.6) and (2.10). As such, in view of non-vanishing 4-brane tensions, it appears on physical grounds that the “brane-box” configuration with walls made up of the four 4-branes is effectively bounded along the $y$-direction by the two 4-branes at $y=0$ and at $y=\pi$, whereas along the $z$-direction it effectively consists of only one non-trivial 4-brane at $z=\pi$. We thus note that the $y$-direction mimics a 6D analogue of the RS1 scenario [1], whereas the $z$-direction is similar to an RS2 scenario [62] which is based on only one brane (with the other one formally taken to be at infinity). The fundamental 6D braneworld model under our consideration may thus be interpreted as a combination of individual RS1 and RS2 scenarios along the two extra dimensions. Moreover, according to the analysis in [61], this interpretation holds for any arbitrary $n$-fold nested warped geometry, where the model is analogous to a $(4+n)$-dimensional RS1 scenario only along the first orbifold, and mimics $(4+n)$-dimensional RS2 scenarios along all the subsequent (higher) orbifolds.
@@ 3 Effective four-dimensional dynamics
Before proceeding directly with these calculations however, it is imperative to highlight how the effective 4D Planck scale ($M_{\rm eff}$) is connected to $M$, i.e. the fundamental Planck scale. This relation sets the strength of the couplings in the effective 4D theory and serves as a bridge between the latter and the fundamental higher-dimensional theory. As shown in [33], the relation between $M_{\rm eff}$ and $M$ in the flat doubly warped braneworld model depends on the moduli $c$ and $k$, and can be expressed as
$M_{\rm eff}^{2}=\dfrac{2M^{4}}{k^{\prime 2}}\cosh(k\pi)\left(1-e^{-2c\pi}\right)\left(1+\dfrac{1}{3}\sinh^{2}(k\pi)\right)\tanh(k\pi)\operatorname{sech}(k\pi)\>,$ (3.1)
where the mass scale $k^{\prime}=\sqrt{-\dfrac{\Lambda_{6}}{40M^{4}}}$ has been defined. With $\Lambda_{6}\sim-M^{6}$, we end up with $k^{\prime}\sim M$. Thus, for dominant warping along either orbifold direction, i.e. with either $c\gtrsim\mathcal{O}(1)$ or $k\gtrsim\mathcal{O}(1)$, it is evident from (3.1) that the effective 4D Planck mass remains very close to the fundamental Planck mass -- a feature similar to the standard 5D scenario [1]. In the generalized non-flat scenario under our consideration, the 4D Planck mass can be obtained in a similar manner, being defined by
[…]
which leads to different relations in the dS and AdS branches according to (2.3) and (2.7). These relations can be analytically simplified to
  • AdS:
    $M_{\rm eff}^{2}=\dfrac{1}{12c_{1}^{2}}\left(\dfrac{M^{4}}{k^{\prime 2}}\right)$$\times\left[\left(1-e^{-2c\pi}\right)c_{1}^{4}+4\pi c\>w_{1}^{2}c_{1}^{2}+\left(e^{2c\pi}-1\right)w_{1}^{4}\right]$ (3.3)
    ×$\times\left[5+\cosh(2k\pi)\right]\tanh(k\pi)\operatorname{sech}(k\pi)\>,$
  • dS:
    $M_{\rm eff}^{2}=\dfrac{1}{12c_{2}^{2}}\left(\dfrac{M^{4}}{k^{\prime 2}}\right)$$\times\left[\left(1-e^{-2c\pi}\right)c_{2}^{4}-4\pi c\>w_{2}^{2}c_{2}^{2}+\left(e^{2c\pi}-1\right)w_{2}^{4}\right]$ (3.4)
    ×$\times\left[5+\cosh(2k\pi)\right]\tanh(k\pi)\operatorname{sech}(k\pi)\>.$
For $w_{1,2}\ll 1$ and $c_{1,2}=1+\sqrt{1\mp w_{1,2}^{2}}$ as defined earlier, it is clear that dominant warping along either orbifold results in $M_{\rm eff}\approx M$, which is similar to the flat case. This proximity of the effective Planck scale to the fundamental scale is, indeed, a generic feature of warped braneworlds. Motivated by this, we do not distinguish explicitly between $M_{\rm eff}$ and $M$ in subsequent calculations.
@@ 3.2 Effective kinetic and curvature terms
with the mass scale $k^{\prime}=\sqrt{-\dfrac{\Lambda_{6}}{40M^{4}}}$ extracted to make the Jordan frame coupling $h(c,k)$ dimensionless. The explicit forms of $h(c,k)$ for the AdS and dS cases, obtained by plugging in the corresponding warp factors in the integral above, are given in (A.1) and (A.2) in Appendix A.
@@ 3.3 Transformation from Jordan to Einstein frame
Before proceeding further, some clarifications regarding the dimensions of the various quantities are in order. In the first integral, $[R]=+2$. In the second integral, the modular potential $V$ is dimensionless by construction. Finally, in the third integral, the non-canonical coefficients are dimensionless as well, rendering each individual kinetic term of dimension $+2$. With $k^{\prime}$ being the derived mass scale, the entire 4D effective action is thus dimensionless as required. Clearly, this action contains a non-minimal coupling of curvature to the moduli fields through $h(c,k)$. This coupling can be made minimal by transforming to the Einstein frame, which is enacted by first assuming
@@
Figure 1: Representative plots of the Einstein frame non-canonical kinetic coupling functions $\hat{\tilde{A}}_{kk}(k)$ on the upper left, $\hat{\tilde{A}}_{cc}(c)$ on the upper right, $\hat{\tilde{A}}_{ck}(c,k)$ at the lower left, and the Einstein frame induced potential $\hat{V}(c,k)$ at the lower right, for AdS geometry with $\omega_{1}=10^{-15}$. Note that all the quantities are dimensionless, and $\hat{V}(c,k)$ has been shown multiplied by an overall factor of $10^{15}$.
@@
Figure 2: Representative plots of the Einstein frame non-canonical kinetic coupling functions $\hat{\tilde{A}}_{kk}(k)$ on the upper left, $\hat{\tilde{A}}_{cc}(c)$ on the upper right, $\hat{\tilde{A}}_{ck}(c,k)$ at the lower left, and the Einstein frame induced potential $\hat{V}(c,k)$ at the lower right, for dS geometry with $\omega_{2}=10^{-3}$. Note that all the quantities are dimensionless, and $\hat{V}(c,k)$ has been shown multiplied by an overall factor of $10^{6}$.
@@
The kinetic coefficients and the induced potential in the Einstein frame have been plotted in Fig. 1 for the AdS case and in Fig. 2 for the dS case. The explicit forms of the AdS kinetic coefficients are given in Appendix A under (A.1), (A.1), and (A.1), and the dS ones under (A.2), (A.2), and (A.2). The corresponding functional forms of $\hat{V}(c,k)$ appear in (A.1) and (A.2). It is interesting to note that $\hat{\tilde{A}}_{cc}$ depends on $c$ only and $\hat{\tilde{A}}_{kk}$ is a function of $k$ alone, while $\hat{\tilde{A}}_{ck}$ is a function of both $c$ and $k$. In this sense, the Einstein frame allows a nice decomposition of the functional dependence of the coupling coefficients based on the kinetic terms they are associated with, which is a consequence of the underlying nested warped geometry and holds good in higher dimensional extensions. Also, $\hat{\tilde{A}}_{kk}(k)$ is independent of $\Omega$, and as such, is identical across the AdS and dS scenarios. These features appear to be generic properties which can be readily generalized to higher-dimensional extensions, as we discuss later in Sec. 5. On the other hand, $\hat{\tilde{A}}_{cc}(c)$ in the dS case changes sign from negative to positive (for increasing $c$) at a non-zero value of $c$, unlike in the AdS case where it approaches zero from the negative side asymptotically. This feature of the dS regime turns out to be crucial from a cosmological point of view, as we shall demonstrate in the upcoming section.
@@
A few words about $\hat{V}(c,k)$ in the de Sitter regime are in order. In the 5D curved scenario, it is known that the induced dS potential may stabilize the extra-dimensional modulus by virtue of its metastable minimum [58, 59]. While these earlier studies have been carried out in terms of the canonically normalized radion field derived from the modulus, the metastable nature of the plateau should reflect in the modular potential as well. In fact, if one proceeds directly in terms of the modulus in the 5D set-up, the shape of the induced potential that one ends up with is identical to a constant $k$-projection of $\hat{V}(c,k)$ from Fig. 2, as expected physically. Such a constant $k$-slice clearly shows a nearly flat plateau in the range $c\in[1,5]$. For a given value of $\omega_{2}$, this range of $c$ where the metastable region occurs in the $c$-direction does not depend on the value of $k$, whose increment only results in an overall reduction in the height of the metastable interval as shown in Fig. 2. So here is the natural question to ask: might it be possible to stabilize the other moduli of multiply warped geometries in a similar fashion? Based on Fig. 2, we find the answer to be negative. In fact, $\hat{V}(c,k)$ shows no local minimum, metastable or not, in the direction of the second modulus $k$, along which it decays monotonically. As we observe later in Sec. 5, this trend continues to hold in nested warped extensions beyond 6D, with the induced potential displaying no local minimum along the higher-dimensional moduli in field space. Thus, in multiply warped extensions of the 5D curved scenario, one indeed appears to need additional phenomenological mechanisms and/or beyond-Einstein gravitational effects to stabilize the higher moduli beyond the first one.
3.4 Field space curvature in Einstein frame
An important theoretical tool for illustrating the deviation from canonicity is the curvature of the so-called field space metric ($G_{AB}$), which consists of the non-canonical kinetic coefficients associated with the individual kinetic terms. Its definition allows expressing the overall kinetic part of the Lagrangian as $\mathcal{L}_{\rm kin}=G_{AB}\partial_{\mu}\phi^{A}\partial^{\mu}\phi^{B}$, where $A$ and $B$ label the individual fields. In our case, the full field space metric may be represented by a symmetric $2\times 2$ matrix, which comprises of the following components:
$G_{11}=\hat{\tilde{A}}_{cc}(c)\>,\>\>G_{22}=\hat{\tilde{A}}_{kk}(k)\>,\>\>G_{12}=G_{21}=\dfrac{1}{2}\hat{\tilde{A}}_{ck}(c,k)\>,$ (3.16)
where the first two terms are the diagonal terms associated respectively with $\partial_{\mu}c\partial^{\mu}c$ and $\partial_{\mu}k\partial^{\mu}k$, and the last term is the off-diagonal term associated with $\partial_{\mu}c\partial^{\mu}k$. Note that all the components are dimensionless.
Using the tensor computation package xCoba (based on the Wolfram suite xAct [63]), the Ricci curvature scalar ($\mathcal{R}_{\rm field}$) associated with this $G_{AB}$, which is dimensionless owing to the dimensionless nature of $c$ and $k$, is obtained in an analytical form as
$\mathcal{R}_{\rm field}=$ $\dfrac{4}{\left(\hat{\tilde{A}}_{ck}^{2}-4\hat{\tilde{A}}_{cc}\hat{\tilde{A}}_{kk}\right)^{2}}\Bigg{[}\hat{\tilde{A}}_{ck}\left(\dfrac{\partial\hat{\tilde{A}}_{cc}}{\partial c}\dfrac{\partial\hat{\tilde{A}}_{kk}}{\partial k}+\dfrac{\partial\hat{\tilde{A}}_{ck}}{\partial c}\dfrac{\partial\hat{\tilde{A}}_{ck}}{\partial k}\right)$
$-2\left(\hat{\tilde{A}}_{kk}\dfrac{\partial\hat{\tilde{A}}_{cc}}{\partial c}\dfrac{\partial\hat{\tilde{A}}_{ck}}{\partial k}+\hat{\tilde{A}}_{cc}\dfrac{\partial\hat{\tilde{A}}_{kk}}{\partial k}\dfrac{\partial\hat{\tilde{A}}_{ck}}{\partial c}\right)-\left(\hat{\tilde{A}}_{ck}^{2}-4\hat{\tilde{A}}_{cc}\hat{\tilde{A}}_{kk}\right)\dfrac{\partial^{2}\hat{\tilde{A}}_{ck}}{\partial c\>\partial k}\Bigg{]}\>.$
We can use (3.4) to evaluate the field space curvature in both the AdS and the dS branches, by making use of the analytical expressions of the non-canonical kinetic coefficients obtained in each branch (viz. Appendix A). In Fig. 3, we present the behaviour of $\mathcal{R}_{\rm field}$ graphically. Owing to different order-of-magnitude variations in the directions of the two moduli fields, it is difficult to visualize the full $\mathcal{R}_{\rm field}$ surface as a function of both $c$ and $k$. Hence, we show one-dimensional profiles of $\mathcal{R}_{\rm field}$ as a function of one of the fields at a time, for a few benchmark parametric values of the other field.
Some interesting features are readily observable in Fig. 3. Firstly, $\mathcal{R}_{\rm field}<0$ throughout the entire range of field values, which is a smoking gun signature of non-canonicity (canonical kinetic terms should yield $\mathcal{R}_{\rm field}=0$ identically). Secondly, in the AdS case, the scenario approaches canonicity monotonically as $c$ increases (i.e. $\mathcal{R}_{\rm field}\to 0$), whereas the non-canonical nature becomes most pronounced for a finite value of $k$ (represented by the location of the trough in the upper-right plot) beyond which non-canonicity decreases monotonically along higher $k$. In the dS case, on the other hand, non-canonical behaviour becomes extremal for a combination of non-zero values of both $c$ and $k$, represented by the troughs along both the $c$ and the $k$ directions in the two plots in the bottom row of Fig. 3. Beyond these values, non-canonicity gradually decreases for higher values of $c$ and $k$, as the scenario asymptotes towards canonicity. Thirdly, while we have used the same range of parametric values for $c$ and $k$ in the AdS and dS cases, the variation of $\mathcal{R}_{\rm field}$ with respect to these values shows opposite trends to some extent across these two regimes. This becomes clear upon comparing the upper-right and the bottom-right plots of Fig. 3, where the variation of $\mathcal{R}_{\rm field}$ as a function of $k$ is shown for different parametric values of $c$. While the minima become shallower upon increasing the value of $c$ in the AdS case, they become deeper for the same in the dS case. Besides, the absolute value of $\mathcal{R}_{\rm field}$ at the minima are, on an average, one order of magnitude smaller in the AdS case than in the dS case for identical ranges of $c$.
Figure 3: Representative plots of the field space Ricci curvature scalar $\mathcal{R}_{\rm field}$ (dimensionless) in the AdS regime (upper row) and in the dS regime (lower row). Due to ease of visualization, one-dimensional profiles are shown as a function of one of the moduli fields at a time, with the variation with respect to the other one shown parametrically. Similar to Figs. 1 and 2, the plots here correspond to $\omega_{1}=10^{-15}$ in the AdS case and $\omega_{2}=10^{-3}$ in the dS case.